# Samson Lasaulce, Giuseppe Valenzise, Zhuo He, and Merouane Debbah

### v1

@@ II A short review of GenAI techniques
GANs [2] generally consist of two neural networks interacting in the form of a zero-sum game, in which the generator $G$ aims to generate new data with the same statistics as the training dataset while the discriminator $D$ focuses on the identification of the true and fake data samples. In RF sensing applications, GANs offer the potential to generate realistic synthetic RF data, enhance signal quality, and enable robust anomaly detection. Their ability to model complex signal environments and facilitate data augmentation paves the way for more resilient and adaptive RF sensing systems.
VAEs learn an encoder and a decoder to embed data into a continuous latent space, where the decoder uses a parameterized likelihood function [3]. This approach reduces overfitting, making VAEs ideal for large datasets like radio map construction. In RF sensing, VAEs significantly enhance performance by denoising data and reconstructing incomplete information. By understanding the structure and distribution of RF signals, VAEs filter out noise and infer missing parts, creating comprehensive datasets even with partial sensor coverage or data loss. These capabilities make RF sensing systems more robust, adaptable, and effective, improving their performance and reliability in various IoT applications.
DMs use a Markov chain to add random noise to data (forward diffusion) and then denoise it to generate desired samples (reverse diffusion) [4]. The forward process transforms the data distribution into a simple prior (e.g., Gaussian noise), and new data is generated by denoising samples from this prior. Diffusion models enhance RF sensing in IoT by iteratively generating and denoising data, producing high-quality synthetic RF data and improving signal clarity. They effectively detect anomalies and capture fine variations in RF environments, and robustly reconstruct incomplete RF data, ensuring comprehensive datasets despite partial sensor coverage. These capabilities improve the precision, reliability, and robustness of RF sensing in IoT applications.
LLMs [5] are pre-trained on massive text datasets to learn contextual dependencies, and have achieved remarkable success in various applications, including question answering, language comprehension, reading comprehension, code creation, and both mathematical and common sense reasoning. In IoT systems, LLMs enhance RF sensing by incorporating natural language processing, enabling sensors to interpret and generate human language for smarter device communication. By involving the natural language modality, LLMs facilitate sophisticated multi-modal data analysis, combining RF data with textual and audio inputs for comprehensive insights. This capability improves anomaly detection, contextual understanding, and decision-making processes, making IoT systems more intelligent and adaptive.
@@
width = colspec = Q[200]Q[362]Q[378], cells = Thistle, valign = m, cells = BlueChalk, valign = m, column2 = Seashell, cell12 = fg=MineShaft, cell13 = fg=MineShaft, cell22 = fg=MineShaft, hlines, vline2-3 = -, Principal issues & Limitation of traditional deep learning techniques Potential of Generative AI
@@ III GenAI empowered uni-modal wireless sensing in IoT
The field of RF sensing involves utilizing radio frequency signals to detect and interpret physical phenomena, offering critical capabilities for applications such as environmental monitoring, health diagnostics, and security systems. Both traditional and modern RF sensing techniques face inherent challenges during the data acquisition and transmission stages. Among these challenges, high costs for data collection, significant interference during data acquisition, and bandwidth limitations for the data transmission are particularly prominent. Additionally, although RF signals are considered crucial information source in many IoT applications, they are extremely sparse due to the difficulty and high costs associated with setting up extensive observation points. To address these issues, we explore the emerging application of GenAI techniques to uni-modality RF sensing, building on their success in the computer vision domain, and provide visions for future directions.
Moreover, despite the advantages of using RF signals in IoT systems, challenges arise due to the rapid increase in IoT devices, which complicates efficient signal acquisition and data collection. Precise synchronization, robust interference handling, and managing large data volumes are particularly challenging. Consequently, RF signals often suffer from missed observations and exhibit a sparse structure, creating obstacles for effective downstream tasks. Generative AI addresses these challenges by using models like GANs, VAEs, and DMs to reconstruct missing data, infer information for localization, and generate representative samples for downstream tasks (see Fig. 2). The advantages of GenAI in the unimodal case can be discussed in three main aspects.
@@ III-A Data imputation
A specific issue within RF sensing is the handling of missing Received Signal Strength (RSS) readings, which can lead to low accuracy and reliability in various applications, such as localization and environment monitoring. In these cases, the missing data needs to be filled in based on the observed data to minimize the impact. In the context of filling missing data, the correlation between the missing data and observed data is crucial. However, this correlation varies and traditional neural networks face challenges in handling this variability, as they often assume fixed spatial relationships or rely on predefined spatial structures. To address this issue, GenAI models, such as Transformers, have a great potential in learning complex dependencies and varying impacts of different observations, in which attention mechanisms can explicitly model the varying impact of each observation. Additionally, GenAI can create embeddings that capture broader context and relationships within the data, which can generate more accurate missing data. For instance, Wang et al.[6] leverages the Bidirectional Encoder Representations from Transformers (BERT) model, known for its strong representation learning ability and generalization capability in natural language processing, to input collected RSS data and predict missing signals in the radio map. In the case of RF sensing, BERT can be adapted to understand the contextual relationships within RSS data, much like it understands word relationships in a sentence. This capability allows BERT to model complex dependencies and patterns in the RSS data, making it highly effective at inferring missing values. By accurately predicting these missing signals, BERT helps to reduce the workload of collecting RSS in the offline phase and provides accurate positioning results in the online phase. This ultimately improves the efficiency, accuracy, and cost-effectiveness of radio map construction.
@@ III-B Super-resolution
While these advancements are promising, there are still significant challenges to address in RF sensing. In particular, in many IoT deployments, the number of sensors is limited due to cost constraints, power requirements, or physical space limitations. This can result in large areas with minimal sensor coverage, leading to extremely sparse data. When there are large missing regions (e.g., 90% missing), GenAI models are more effective due to their ability to generate data that is not directly present in the input, enabling to infer missing data under large missing rates and complicated loss patterns [7]. Among them, DMs show its biggest potential for super-resolution data generation from extremely sparse latent spaces. DMs generate images through a process of gradually reducing noise, modeling the distribution of possible images at each step. This makes them effective at dealing with incomplete input. Despite the efficiency of DMs, they still bring the problem of low authenticity. This problem could be addressed by the idea of guided diffusion models, which utilize additional information, such as labels, text descriptions, or other types of metadata to guide the model for a more expected output. Building on this concept, LLMs can be employed to create customized guidance for a wide range of sensing tasks, such as tailoring the guiding prompts for radio map or any other patterns.
@@ III-C Synthetic data generation
In IoT scenarios, data collection is challenging due to logistical, financial, and privacy constraints, as well as the variability and complexity of real-world environments. By leveraging AI models to generate high-fidelity synthetic data, it is possible to augment limited real-world data, enhancing the training and performance of RF signal processing algorithms. GenAI can improve model robustness by generating synthetic data from learned distributions, introducing controlled variability, and helping models generalize to unseen conditions. Specifically, GenAI can create data tailored to specific scenarios, such as varying environmental conditions, interference patterns, and device configurations, simulating diverse settings like urban environments, indoor spaces, and rural areas. In [8], a generative adversarial networks for RSS data augmentation is proposed to generate synthetic RSS data based on a small set of real collected labeled data, which achieved to efficiency improving localization accuracy. The inherent statistical character and use case examples in various domains illustrate that GenAI play a crucial role in addressing the data collection issues by generating synthetic data that enhances the training and performance of RF sensing models in IoT scenarios.
@@ IV-A Cross-modal RF sensing with GenAI
In RF sensing scenarios, during the training phase, models may be trained using simulated RF data or data collected under controlled conditions where signals are easy to observe and interference are controlled. However, in real-world deployment, the actual RF signals might be sparse, noisy, or even unavailable due to environmental constraints or hardware limitations. GenAI-empowered cross-modal RF sensing techniques address the problems in various methodologies.
LLMs provide great potential for cross-modal learning with their strong capacity to enhance the generalizability of multimodal systems. A first strategy is to train universal models on vast datasets across numerous modalities, to encode these modalities into embeddings. More specifically, multi-modal LLMs can help to generate embeddings that effectively represent the underlying information from different modalities, therefor helping to enhance the generalizability of the universal models. The second direction, From a more precise perspective with LLM models, is to convert every modality into text modality and build the model based on text prompts. It has the potential to provide robust system for cross-modal sensing when one of the modalities is missing, as text can serve as a unified semantic space, leveraging the extensive zero-shot prediction capabilities of LLMs. For instance, [9] leverages LLMs and foundational models to create a unified semantic space using text. This approach shows the capacity of LLMs to enhance the generalizability of multimodal systems, allowing them to perform zero-shot predictions across different modality combinations during testing.
@@
VAEs provide great potential to learn to encode RF signals into a latent space and decode them into another modality. More specifically, based on the probabilistic nature of VAEs, they have the capacity to adopts a hierarchical approach to encode data into a distribution over latent variables, and generate diverse outputs in cross-modal tasks. VAEs can learn both joint and modality-specific distributions for different data types (e.g., RF signals and images), enabling effective cross-modality inference. This allows the model to leverage the strengths of each modality and make accurate predictions even when some data is missing or incomplete, by inferring the missing parts from the shared latent space. For instance, [10] introduced a Multimodal Hierarchical Variational Auto-encoder (MHVAE), where the single modality nature of VAEs is extended to a multimodal hierarchical setting, enabling the learning of modality-specific distributions and a joint-modality distribution for cross-modality inference. It showcasts the effectiveness of VAEs in learning rich multimodal representation and cross-modal generation.
DMs also offers a effective solution for synthesizing cross-modality data by leveraging existing datasets. Based on hierarchical statistical structure and iterative denoising process, DMs show its capacity for tasks requiring high-quality and detailed cross-modal outputs, such as text-to-image generation. What’s more, DMs can effectively capture and model complex relationships between different modalities, making them suitable for sophisticated cross-modal tasks, such as text-to-image generation. For instance, in [11], an effective method for synthesizing RF sensing data using cross-modal diffusion models to enhance the generalization capability of millimeter-wave (mmWave) sensing systems is proposed, demonstrating a remarkable enhancement in sensing and generalization capabilities across two distinct sensing tasks.
Besides, GenAI has the potential to revolutionize RF sensing by enabling personalized and context-aware applications that enhance user experiences and improve the efficiency of various systems. For instance, conditional GenAI techniques can be trained to generate data conditioned on specific inputs, such as text descriptions and contextual data like location or user preferences, therefor integrating contextual information to generate relevant and personalized outputs.
@@ IV-B Multi-modal fusion with GenAI
Figure 4: Illustration of the RF-to-image Generative AI model utilizing a text-prompt conditioned Transformer, and the reconstructed building obstacles image from real measurement. Initially, a Convolutional Neural Network (CNN) generates rough images from extremely sparse receive power samples of RF signals, which are blurred and noisy. These blurred images are then refined by GPT4o with a text prompt, enhancing image quality by removing noise and sharpening details. The figure highlights the significant improvements achieved through generative AI, demonstrating its capability to generate images with low distortion and high perceptual quality, thereby showcasing its advantages in RF sensing applications.
Enhancing RF Sensing with Language Descriptions: 1) Consider a scenario where a network of sensors is deployed in an environment for monitoring purposes. While the primary RF sensors capture signals that provide spatial and movement RF data, supplementary sensors record descriptive information about the environment. These descriptions might include details such as ”three people standing near the entrance,” ”a moving vehicle is approaching from the west,” or ”there is an obstruction in the path.” Transmitting these descriptions instead of raw data minimizes the bandwidth requirements and simplifies data collection. 2) Using LLMs, these language descriptions can be fused with the RF signal data to enhance downstream tasks such as localization, object detection, and activity recognition. LLMs can process and integrate the descriptive information, correlating it with the RF data to create a more complete and accurate representation of the monitored environment. Moreover, the flexibility of LLMs allows them to interpret the descriptive data, making them adapted to scenarios where traditional sensing methods might struggle. For instance, in dynamic environments where RF signals might be highly variable, the additional context provided by language descriptions can significantly improve the reliability and accuracy of the sensing system. A hybrid End-to-End learning framework for autonomous driving by combining basic driving imitation learning with LLMs based on multi-modality prompt tokens is proposed in [13], where end-to-end integration of visual and LiDAR sensory are input into learnable multi-modality tokens, and a hybrid setting is explored which use LLMs to help the driving model correct mistakes and complicated scenarios.
@@ IV-C Case study: GenAI models for RF imaging
GenAI models have been widely used to generate data cross different modalities. For example, DALLE-2 uses diffusion model or auto-regressive model to generate image embedding from text embedding, plus contrastive learning to align the embeddings on pair-wise data. Moreover, causal transformer and U-Net are shown as effective architecture to implement the diffusion prior and diffusion generator. Inspired by this, we propose a RF to image GenAI model based text-prompt conditioned Transformer model. As illustrated in 4, the figure depicts the structure of our proposed method. Initially, a Convolutional Neural Network (CNN) is trained to generate images from extremely sparse receive power samples of RF signals, producing a rough representation often characterized by blur and noise. These blurred images are then fed into a Transformer model conditioned by a text prompt, which further enhances the image quality by leveraging its learned data distribution to remove noise and sharpen details. The comparison between the intermediate blurred images and the final output highlights the significant improvements achieved through generative AI. While traditional neural networks struggle to generate clear images from sparse data due to their limited capacity to learn exact data distributions, the Transformer model outperforms by providing high-quality reconstructions. This case study can show the capability of GenAI to generate images with both low distortion and high perceptual quality, highlighting its significant advantage in RF sensing applications.

### v2

# Generative AI for RF Sensing in IoT systems

@@ II A short review of GenAI techniques
GANs consist of two neural networks engaging in a zero-sum game: the generator $G$ aims to create new data that matches the statistics of the training dataset, while the discriminato $D$ identifies true and fake data samples. In RF sensing applications, GANs can generate realistic synthetic RF data, enhance signal quality, and enable robust anomaly detection. Their capability to model complex signal environments and facilitate data augmentation contributes to more resilient and adaptive RF sensing systems.
VAEs learn an encoder and decoder to embed data into a continuous latent space, using a parameterized likelihood function that reduces overfitting—making them ideal for large datasets like radio map construction. In RF sensing, VAEs enhance performance by denoising data and reconstructing incomplete information. By understanding the structure and distribution of RF signals, they filter out noise and infer missing parts, creating comprehensive datasets despite partial sensor coverage or data loss. These capabilities improve the robustness, adaptability, and effectiveness of RF sensing systems in various IoT applications.
DMs utilize a Markov chain to add random noise to data (forward diffusion) and then denoise it to generate desired samples (reverse diffusion). This process transforms the data distribution into a simple prior (e.g., Gaussian noise), allowing new data to be generated through denoising. DMs enhance RF sensing in IoT by iteratively generating and denoising data, producing high-quality synthetic RF data and improving signal clarity. They effectively detect anomalies, capture fine variations in RF environments, and robustly reconstruct incomplete RF data, ensuring comprehensive datasets despite partial sensor coverage. These capabilities enhance the precision, reliability, and robustness of RF sensing in IoT applications.
LLMs are pre-trained on vast text datasets to learn contextual dependencies, achieving significant success in applications like question answering, language comprehension, code creation, and reasoning. In IoT systems, LLMs enhance RF sensing by incorporating natural language processing, enabling sensors to interpret and generate human language for smarter device communication. By integrating the natural language modality, LLMs facilitate sophisticated multi-modal data analysis, combining RF data with textual and audio inputs for comprehensive insights. This capability improves anomaly detection, contextual understanding, and decision-making, making IoT systems more intelligent and adaptive.
@@
width = colspec = Q[200]Q[362]Q[378], cells = Thistle, valign = m, cells = BlueChalk, valign = m, column2 = Seashell, cell12 = fg=MineShaft, cell13 = fg=MineShaft, cell22 = fg=MineShaft, hlines, vline2-3 = -, issues & Limitation of traditional deep learning techniques Potential of Generative AI
@@ III GenAI empowered uni-modal wireless sensing in IoT
RF sensing utilizes radio frequency signals to detect and interpret physical phenomena, providing essential capabilities for applications like environmental monitoring, health diagnostics, and security systems. Both traditional and modern RF sensing techniques face challenges during data acquisition and transmission, including high costs, significant interference, and bandwidth limitations. Although RF signals are vital in many IoT applications, their sparse nature results from the difficulty and expense of establishing extensive observation points.
To address these challenges, we explore the application of Generative AI techniques in uni-modality RF sensing, inspired by their success in computer vision. As the number of IoT devices increases, efficient signal acquisition and data collection become more complex, leading to issues with synchronization, interference handling, and large data volumes. Consequently, RF signals often suffer from missed observations and sparse structures, hindering effective downstream tasks. Generative AI models, such as GANs, VAEs, and diffusion models, can help reconstruct missing data, infer information for localization, and generate representative samples for downstream tasks (see Fig. 2). The advantages of Generative AI in this context can be categorized into three main aspects.
@@ III-A Data imputation
A key issue in RF sensing is managing missing Received Signal Strength (RSS) readings, which can reduce accuracy and reliability in applications like localization and environment monitoring. To mitigate this, missing data must be filled in based on observed data, with the correlation between the two being crucial. However, this correlation can vary, and traditional neural networks struggle to accommodate this variability due to their reliance on fixed spatial relationships. Generative AI models like Transformers are well-suited to address the challenge of missing RSS readings. They learn complex dependencies and the varying impacts of different observations through attention mechanisms that model the significance of each data point. Additionally, GenAI creates embeddings that capture broader context and relationships, improving predictions for missing data. For example, Wang et al. [2] utilize the Bidirectional Encoder Representations from Transformers (BERT) model to effectively infer missing values by understanding the contextual relationships within RSS data. More broadly, the imputation technique using GenAI is particularly valuable for IoT applications with incomplete RF measurements. In smart city infrastructure, for example, environmental factors or structural obstructions can cause intermittent signal loss. GenAI can fill in missing RF data, ensuring consistent monitoring for applications like traffic flow analysis. This approach also benefits intelligent transportation systems, where IoT-based tracking encounters variable RF conditions across different regions. By learning from historical patterns, GenAI can deliver reliable data streams, supporting accurate and uninterrupted monitoring in complex IoT environments.
@@ III-B Super-resolution
While these advancements are promising, there are still significant challenges to address in RF sensing. In particular, in many IoT deployments, the number of sensors is limited due to cost constraints, power requirements, or physical space limitations. This can result in large areas with minimal sensor coverage, leading to extremely sparse data. When there are large missing regions (e.g., 90% missing), GenAI models are more effective due to their ability to generate data that is not directly present in the input, enabling to infer missing data under large missing rates and complicated loss patterns [3]. Among them, DMs have significant potential for super-resolution data generation from sparse latent spaces, especially in complex IoT applications with limited sensor density. In large-scale industrial monitoring, such as for bridges or pipelines, DMs can convert sparse RF input data into detailed condition maps, enhancing structural health monitoring while minimizing sensor deployment. To improve precision, conditioning diffusion models can integrate contextual data, like structural specifications for industrial monitoring or traffic patterns for urban planning, refining the generated data to meet specific needs. Additionally, Large Language Models (LLMs) can be used to create customized guidance for various sensing tasks, such as tailoring prompts for radio maps.
@@ III-C Synthetic data generation
In IoT scenarios, data collection faces logistical, financial, and privacy constraints, along with the complexity of real-world environments. By leveraging AI models to generate high-fidelity synthetic data, it’s possible to augment limited real-world data, enhancing the training and performance of RF signal processing algorithms. Generative AI can improve model robustness by creating synthetic data from learned distributions, introducing controlled variability, and enabling generalization to unseen conditions. Specifically, GenAI can produce data tailored to scenarios with varying environmental conditions, interference patterns, and device configurations, simulating diverse settings like urban, indoor, and rural areas. For instance, Njima et al. [4] proposed a generative adversarial network for RSS data augmentation, successfully generating synthetic RSS data from a small set of real labeled data, which improved localization accuracy. Synthetic data generation is especially important for applications where real-world data is scarce or costly to obtain. In industrial IoT, synthetic data can simulate harsh environmental conditions and operational variability, allowing models to better predict equipment failures and optimize maintenance schedules. Similarly, in healthcare IoT, where privacy constraints limit data availability, synthetic RF data can enhance model training for indoor localization and patient monitoring systems, supporting accurate and reliable performance across diverse healthcare settings.
@@ IV-A Cross-modal RF sensing with GenAI
In RF sensing, models are often trained with simulated data or in controlled conditions where signals are clear. However, real-world deployment may involve sparse, noisy, or unavailable RF signals due to environmental constraints or hardware limitations. GenAI-empowered cross-modal RF sensing techniques address these problems through various methodologies.
LLMs offer significant potential for cross-modal learning, enhancing the generalizability of multimodal systems. One strategy involves training universal models on large datasets across various modalities to encode these modalities into embeddings. Multi-modal LLMs generate embeddings that effectively represent underlying information, improving the generalizability of universal models. Another approach involves converting all modalities into text, creating a model based on text prompts. This method provides a robust system for cross-modal sensing when a modality is missing, as text acts as a unified semantic space, utilizing LLMs’ zero-shot prediction capabilities. For example, [5] demonstrates how LLMs and foundational models can create a unified semantic space using text, showcasing their ability to perform zero-shot predictions across different modality combinations during testing.
@@
VAEs provide great potential to learn to encode RF signals into a latent space and decode them into another modality. Their probabilistic framework allows for a hierarchical approach to encode data as distributions over latent variables, facilitating diverse outputs in cross-modal tasks. VAEs can learn both joint and modality-specific distributions, enabling effective cross-modality inference, which allows for accurate predictions even with missing or incomplete data by inferring from the shared latent space. For example, [6] introduced a Multimodal Hierarchical Variational Autoencoder (MHVAE), extending the single-modality nature of VAEs to a multimodal hierarchical setting. This approach enables the learning of both modality-specific and joint distributions, demonstrating VAEs’ effectiveness in rich multimodal representation and cross-modal generation.
DMs DMs also provide an effective solution for synthesizing cross-modal data by leveraging existing datasets. Their hierarchical statistical structure and iterative denoising process enable high-quality, detailed outputs for tasks like text-to-image generation. DMs adeptly capture and model complex relationships between modalities, making them suitable for advanced cross-modal applications. For instance, [7] proposes a method for synthesizing RF sensing data using cross-modal diffusion models, enhancing the generalization capability of millimeter-wave (mmWave) sensing systems and demonstrating significant improvements in sensing and generalization across two distinct tasks.
Cross-modal generation utilizes GenAI to enhance RF data by learning correlations with more costly or difficult-to-obtain modalities like images and audio. This is crucial in IoT scenarios where RF signals are plentiful but complementary data is scarce. For example, in industrial IoT, RF signals monitor equipment, while inferred thermal data enhances fault detection without the need for expensive thermal sensors. In remote environmental monitoring, RF data provides continuous feedback, supplemented by cross-modal generation to replace costly satellite imagery. This approach allows IoT systems to leverage cost-effective RF data to fill gaps typically covered by more resource-intensive modalities.
@@ IV-B Multi-modal fusion with GenAI
Figure 4: Illustration of the RF sensing technique for building barrier detection based on a Generative AI model utilizing a Segmentation Diffusion model (SegDiff). Initially, a Convolutional Neural Network (CNN) generates a blurred building barrier map from extremely sparse RF signal receive power samples. These blurred barrier maps are then input into the segmentation diffusion model as a condition, which enhances the quality of the detected barrier map by reducing noise and sharpening details. For performance evaluation, the test dataset is divided into 10 classes, with each class representing data containing a specific number of barriers; images with more than 10 barriers are categorized into class 10. The final performance is assessed using the mean Intersection over Union (mean IoU) metric, reflecting the average IoU value across all data within each class. This figure highlights the significant improvements achieved through GenAI, particularly in scenarios with fewer building barriers in the map. The comparison demonstrates the model’s ability to detect barrier maps with high accuracy and perceptual quality, showcasing its advantages in RF sensing-based IoT applications.
@@
Enhancing RF Sensing with Language Descriptions: 1) Consider a scenario where a network of sensors is deployed in an environment for monitoring purposes. While the primary RF sensors capture signals that provide spatial and movement RF data, supplementary sensors record descriptive information about the environment. These descriptions might include details such as “three people standing near the entrance,” “a moving vehicle is approaching from the west,” or “there is an obstruction in the path.” Transmitting these descriptions instead of raw data minimizes the bandwidth requirements and simplifies data collection. 2) Using LLMs, these language descriptions can be fused with the RF signal data to enhance downstream tasks such as localization, object detection, and activity recognition. LLMs can process and integrate the descriptive information, correlating it with the RF data to create a more complete and accurate representation of the monitored environment. Moreover, the flexibility of LLMs allows them to interpret the descriptive data, making them adapted to scenarios where traditional sensing methods might struggle. For instance, in dynamic environments where RF signals might be highly variable, the additional context provided by language descriptions can significantly improve the reliability and accuracy of the sensing system. A hybrid End-to-End learning framework for autonomous driving by combining basic driving imitation learning with LLMs based on multi-modality prompt tokens is proposed in [9], where end-to-end integration of visual and LiDAR sensory are input into learnable multi-modality tokens, and a hybrid setting is explored which use LLMs to help the driving model correct mistakes and complicated scenarios.
@@ IV-C Case study: GenAI models for RF imaging
To validate the effectiveness of our Generative AI-empowered RF sensing in IoT systems, we focus on the outdoor environment reconstruction problem using a denoising diffusion model, specifically applying RF sensing to detect barriers in smart city applications. Outdoor environment reconstruction involves creating digital representations of outdoor spaces, including terrain, buildings, and infrastructure. This process is crucial for urban planning, augmented reality, simultaneous localization and mapping, and industrial network optimization, as it supports infrastructure deployment, signal propagation analysis, and resource allocation. In our approach, user equipment (UE) and base stations (BS) sample RF signals at extremely sparse locations. These sparsely sampled signals are used for outdoor environment reconstruction, but achieving high-quality barrier detection is challenging due to the limited sample density, which can impact the precision and reliability of the detection process.
GenAI models have been widely used to generate data cross different modalities. For example, causal transformer and U-Net are shown as effective architecture to implement the diffusion prior and diffusion generator. Inspired by these advancements, we propose an RF sensing-building barrier detection method based on Segmentation Diffusion GenAI model (SegDiff). As illustrated in Figure 4, the figure highlights the structure of our proposed method, showcasing the advantages of the Generative AI approach over traditional CNN methods in terms of visual quality and detection accuracy, as measured by Intersection over Union (IoU) performance. Initially, a Convolutional Neural Network (CNN) is trained to generate building barrier maps from extremely sparse receive power samples [10], producing a rough representation often characterized by blur and noise. These blurred building barrier map are then fed as the condition of a Segmentation Diffusion model [11], which enhances the quality of estimated barrier map by leveraging its learned data distribution to remove noise and sharpen details.
For performance evaluation, we categorize the test RF sensing dataset into 10 classes, with each class representing data containing a specific number of building barriers. Data with more than 10 barriers are classified as class 10. We assess overall detection accuracy using the mean Intersection over Union (mean IoU) metric, which represents the average IoU value across all data within each class. The comparison of barrier detection performance between the CNN method and the SegDiff method demonstrates significant improvements achieved through the use of GenAI. While traditional neural networks struggle to generate clear barrier map from sparse RF sensing data due to their limited capacity to learn exact data distributions, GenAI model outperforms by providing high-quality sensing results.